# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ 5 Experiments
5.1 Settings
Models, Datasets, and Evaluation. Following prior work on Bayesian low-rank adaptation (Yang et al., 2023; Wang et al., 2023, 2024d), we use llama-2-7b-hf as our primary LLM backbone. We provide additional results on recent LLM architectures in Sec. 5.4, including Meta-Llama-3-8B, Meta-Llama-3.1-8B, and Mistral-7B-v0.3 from the Llama (Dubey et al., 2024) and Mistral (Jiang et al., 2023) families.
@@ 5.1 Settings
For in-distribution experiments, we evaluate model performance on six commonsense reasoning tasks: Winogrande-Small (WG-S) and Winogrande-Medium (WG-M(Sakaguchi et al., 2021), ARC-Challenge (ARC-C) and ARC-Easy (ARC-E(Clark et al., 2018), Open Book Question Answering (OBQA(Mihaylov et al., 2018), and BoolQ (Clark et al., 2019). Furthermore, we use models fine-tuned on OBQA (Mihaylov et al., 2018) to evaluate their generalization ability on out-of-distribution datasets: college-level chemistry (Chem) and physics (Phy) subsets of MMLU (Hendrycks et al., 2021). Label spaces and prompt templates are detailed in Appendix B.1.
@@
To assess uncertainty estimation, we measure Expected Calibration Error (ECE (Naeini et al., 2015)) and Negative Log-Likelihood (NLL) on the test dataset. We also report Accuracy (ACC) to ensure models maintain strong performance while improving calibration. Additional evaluation details are provided in Appendix B.2.
@@
Table 2: Performance of TFB with different posterior families applied to the mean of BLoB on Llama2-7B pre-trained weights. FR: Full-rank isotropic Gaussian noises are applied to $\Delta{\bm{W}}$; C-STD: Standard deviation matrix ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$ is constant. The evaluation protocol strictly follows Table 1. “Rk.”: Average ranking of each method when compared to all other approaches in the study. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively (only for TFB variants).
@@
ACC ($\uparrow$) BLoB-Mean 72.15$\pm$0.17 69.56$\pm$0.91 86.31$\pm$0.37 75.47$\pm$1.36 82.53$\pm$0.74 86.69$\pm$0.08 - 69.93$\pm$1.20 76.88$\pm$0.41 41.67$\pm$2.25 31.94$\pm$1.77
 + TFB (FR) 67.96$\pm$1.79 71.40$\pm$1.73 85.97$\pm$0.57 73.87$\pm$0.28 81.67$\pm$1.17 87.10$\pm$0.19 1.67 67.23$\pm$0.68 77.58$\pm$0.57 41.33$\pm$2.31 33.33$\pm$5.03
 + TFB (C-STD) 67.76$\pm$2.18 67.79$\pm$1.86 85.50$\pm$0.89 72.92$\pm$0.24 82.53$\pm$0.70 86.19$\pm$0.65 2.83 69.71$\pm$0.98 80.22$\pm$0.20 45.67$\pm$0.58 32.00$\pm$2.65
 + TFB (Ours) 69.94$\pm$1.68 70.72$\pm$2.25 86.74$\pm$0.97 73.13$\pm$2.38 83.13$\pm$0.76 86.36$\pm$0.26 1.50 70.38$\pm$1.03 80.16$\pm$0.71 42.67$\pm$1.15 30.67$\pm$1.53
ECE ($\downarrow$) BLoB-Mean 18.63$\pm$0.31 22.51$\pm$0.93 9.64$\pm$0.60 11.58$\pm$1.24 8.65$\pm$0.98 2.88$\pm$0.07 - 14.00$\pm$1.02 10.70$\pm$0.39 15.05$\pm$0.77 22.90$\pm$2.27
 + TFB (FR) 15.15$\pm$1.63 19.32$\pm$1.55 3.55$\pm$0.07 6.24$\pm$0.55 4.33$\pm$1.60 2.12$\pm$0.36 2.50 7.24$\pm$1.88 4.21$\pm$1.30 10.70$\pm$1.51 13.67$\pm$0.83
 + TFB (C-STD) 11.54$\pm$2.70 7.91$\pm$1.99 2.69$\pm$0.45 4.17$\pm$0.83 4.31$\pm$0.86 5.72$\pm$0.18 1.83 9.22$\pm$0.56 5.29$\pm$0.54 9.59$\pm$2.38 18.51$\pm$1.53
 + TFB (Ours) 6.33$\pm$1.04 5.77$\pm$0.32 3.03$\pm$0.43 4.07$\pm$1.65 5.94$\pm$0.46 5.37$\pm$0.44 1.67 12.28$\pm$1.24 8.07$\pm$1.01 12.36$\pm$1.73 22.02$\pm$0.30
NLL ($\downarrow$) BLoB-Mean 0.79$\pm$0.01 1.27$\pm$0.02 0.57$\pm$0.03 0.60$\pm$0.03 0.56$\pm$0.00 0.32$\pm$0.01 - 0.89$\pm$0.02 0.67$\pm$0.02 1.44$\pm$0.00 1.53$\pm$0.02
 + TFB (FR) 0.73$\pm$0.08 1.13$\pm$0.07 0.40$\pm$0.01 0.54$\pm$0.01 0.49$\pm$0.01 0.31$\pm$0.00 2.00 0.82$\pm$0.01 0.59$\pm$0.01 1.32$\pm$0.02 1.46$\pm$0.03
 + TFB (C-STD) 0.67$\pm$0.06 0.82$\pm$0.03 0.39$\pm$0.02 0.54$\pm$0.01 0.48$\pm$0.02 0.33$\pm$0.01 1.33 0.79$\pm$0.02 0.57$\pm$0.01 1.32$\pm$0.02 1.44$\pm$0.06
 + TFB (Ours) 0.62$\pm$0.03 0.86$\pm$0.01 0.42$\pm$0.03 0.56$\pm$0.03 0.50$\pm$0.01 0.34$\pm$0.00 2.33 0.84$\pm$0.03 0.61$\pm$0.01 1.35$\pm$0.01 1.46$\pm$0.06
Baselines. We compare TFB with state-of-the-art uncertainty estimation methods for LoRA-adapted LLMs, including ensemble-based method: Deep Ensemble (ENS(Lakshminarayanan et al., 2017; Balabanov & Linander, 2024; Wang et al., 2023), variational inference methods: Monte-Carlo Dropout (MCD(Gal & Ghahramani, 2016), Bayes By Backprop (BBB(Blundell et al., 2015), Bayesian LoRA by Backprop (BLoB(Wang et al., 2024d), and post-training method: Laplace-LoRA (LAP(Yang et al., 2023). For reference, we also include two standard PEFT baselines: Maximum Likelihood Estimation (MLE(Hu et al., 2022) and Maximum A Posteriori (MAP). All baselines are implemented following the protocols established in BLoB, with details provided in Appendix B.
TFB Implementation.TFB requires no training and can be directly applied to trained LoRA adapters. We evaluate it on three off-the-shelf LoRA checkpoints: MLE, MAP, and the mean component of BLoB (obtained by discarding BLoB’s standard deviation matrix ${\bm{\Omega}}$). For implementation, we use accuracy as the evaluation metric $l$ and randomly sample a subset of the training data as the anchor dataset ${\mathcal{D}}$ (for the size of each dataset, refer to Appendix B.1). The performance change tolerance $\epsilon$ is set adaptively to either 1% or 0.5%, depending on the checkpoint’s overfitting characteristics. To determine the optimal $\sigma_{q}^{*}$, we conduct parallel experiments with eight values of $\sigma_{q}\in$ [0.01, 0.015, 0.02, 0.025, 0.03, 0.035, 0.04, 0.05] using a single random seed. We construct an approximate function $\widehat{\sigma}_{q}(p)$ through piecewise linear interpolation of the observed performance and estimate $\sigma_{q}^{*}\approx\widehat{\sigma}_{q}(p_{0}-\epsilon)$. Similar to other baseline methods, the final results of TFB are reported as averages across three random seeds using $\sigma_{q}^{*}$.
Table 1 presents comprehensive performance metrics (ACC, ECE, and NLL) for various methods applied to LoRA on Llama2-7B pre-trained weights. Due to space limitations, we report the full empirical results of TFB variants (with different $\epsilon$s) in Appendix C.1.
In-Distribution Results. Looking at the IND Datasets results, several key patterns emerge. The MLE baseline shows relatively strong accuracy but suffers from high ECE values (e.g., 29.83 on WG-S), indicating significant overconfidence. This aligns with the common challenge of LLM overconfidence during conventional fine-tuning.
TFB applied to BLoB-Mean demonstrates strong overall performance across the IND datasets, achieving the highest accuracy on several datasets (69.94 on WG-S, 70.72 on ARC-C, and 86.74 on ARC-E). More importantly, it achieves this while maintaining lower ECE values compared to methods like MCD and ENS, suggesting better calibrated predictions. The method also shows strong NLL performance, with values consistently among the lowest across datasets (0.62 for WG-S, 0.86 for ARC-C).
In summary, TFB consistently enhances the performance of baseline methods (MLE, MAP, and BLoB-Mean) across different evaluation scenarios, with notable improvements in both accuracy and calibration metrics. The improvements are particularly evident in the significant ECE reductions (e.g., from 29.83 to 16.26 for MLE on WG-S) while maintaining or improving accuracy, with the most substantial gains observed when TFB is combined with BLoB-Mean, achieving both the highest accuracy and lowest ECE values across most datasets.
Out-of-Distribution Results. The OOD evaluation reveals interesting patterns across both smaller and larger distribution shifts. For smaller shifts (ARC-C and ARC-E), BLoB-Mean with TFB maintains strong performance, achieving 70.38 and 80.16 accuracy respectively, while keeping ECE values low (12.28 and 8.07). This suggests robust generalization under moderate distribution shifts.
For larger shifts (Chem and Phy datasets), we see a more significant performance degradation across all methods, as expected. However, BLoB-Mean with TFB still maintains competitive performance, achieving 42.67 accuracy on Chem and 30.67 on Phy, while maintaining reasonable calibration metrics. The method’s NLL values (1.35 and 1.46 respectively) remain competitive with other approaches, indicating relatively well-calibrated uncertainty estimates even under substantial distribution shifts.
Notable is the consistently strong performance of the BLoB variants (both w/ and w/o TFB) across different metrics and datasets, suggesting that this approach offers a robust framework for both in-distribution and out-of-distribution scenarios. The results demonstrate that the method successfully balances the trade-off between accuracy and calibration, particularly evident in the OOD scenarios where maintaining both aspects becomes more challenging.
5.3 TFB Beyond Low-Rank Isotropic Gaussians
@@
In this section, we extend TFB to other families of Gaussians for modeling the variational distributions of ${\bm{W}}$. Specifically, we consider: (i) Full-Rank Isotropic Gaussian (FR, ${\bm{\Sigma}}_{q}=\sigma_{q}^{2}{\bm{I}}$), and (ii) Constant Low-Rank Standard Deviation (C-STD, ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$). For fair comparison, we adopt the same optimal $\sigma_{q}^{*}$ search protocol of TFB as decribed in Sec. 5.1. The performances of TFB and its variants applied to the mean of BLoB are reported in Table 2.
The experimental results show that TFB outperforms both the baselines FR and C-STD across multiple metrics on in-distribution datasets, with notable improvements in accuracy (e.g., OBQA: 83.13) and calibration (ECE reduced by up to 16.74). While the baseline methods show better NLL scores, these improvements come at the cost of significantly degraded overall performance, making them impractical for real-world applications. While it maintains strong performance on datasets with smaller distributional shifts, its advantages diminish on datasets with larger shifts in Physics and Chemistry domains.
Advantages of TFB.TFB offers additional advantages over the other families of variational distributions discussed above. It is computationally more efficient than FR with noise complexity of $O(rn)$ versus $O(mn)$. Furthermore, unlike C-STD whose variational distributions vary with different but equivalent LoRA matrix pairs and can potentially break the conditions in Theorem 4.2 (Appendix C.2), TFB produces consistent Bayesianization for all equivalent LoRAs satisfying ${\bm{B}}{\bm{A}}=\Delta{\bm{W}}$.
Table 3: Performance of different LLM backbones on the combined dataset of six commonsense reasoning tasks. Boldface denotes the best performance.
@@
5.4 TFB Beyond the Llama2-7B Backbone
We conduct comprehensive experiments across multiple LLM backbones to validate our approach. Our experiments span several models from the Llama architecture (Touvron et al., 2023b; Dubey et al., 2024), including llama-2-7b-hf, Meta-Llama-3-8B, and Meta-Llama-3.1-8B. While we initially considered Llama-3.2-1B, we ultimately discarded its results due to poor adaptation performance with the smaller model architecture. We also extend our analysis to include Mistral-7B-v0.3 (Jiang et al., 2023).
Following commonsense-170k (Hu et al., 2023; Wang et al., 2024c), we combine the 6 reasoning sub-tasks from the main experiments with one shared label space (Combined) and train the base LoRA adapters with MLE. Break-down statistics of each sub-dataset are available in Appendix C.3. For TFB, we set the variational posterior’s standard deviation to $\sigma_{q}=$1e-3 for all three Llama models, while using a lower value of $\sigma_{q}=$2e-3 for Mistral-7B-v0.3.
The results shown in Table 3 demonstrate TFB’s effectiveness: it dramatically reduces ECE across all models (e.g., from 4.74 to 1.05 for Llama3.1-8B) while maintaining strong accuracy and NLL scores with minimal degradation.

### v2

@@ 5.1 Settings
ACC ($\uparrow$) MCD 78.03$\pm$0.61 81.64$\pm$1.79 91.37$\pm$0.38 83.18$\pm$0.84 87.20$\pm$1.02 89.93$\pm$0.16 81.42$\pm$1.38 87.27$\pm$0.84 47.92$\pm$2.25 46.53$\pm$0.49
ENS 78.82$\pm$0.52 82.55$\pm$0.42 91.84$\pm$0.36 83.99$\pm$0.74 87.37$\pm$0.67 90.50$\pm$0.14 79.62$\pm$0.57 86.56$\pm$0.60 49.65$\pm$3.22 44.44$\pm$1.96
LAP BP 76.05$\pm$0.92 79.95$\pm$0.42 90.73$\pm$0.08 82.83$\pm$0.85 87.90$\pm$0.20 89.36$\pm$0.52 81.08$\pm$1.20 87.21$\pm$1.20 48.26$\pm$3.93 46.18$\pm$1.30
BLoB 76.45$\pm$0.37 82.32$\pm$1.15 91.14$\pm$0.54 82.01$\pm$0.56 87.57$\pm$0.21 89.65$\pm$0.15 79.75$\pm$0.43 87.13$\pm$0.00 42.71$\pm$3.71 44.79$\pm$6.64
MLE - 77.87$\pm$0.54 81.08$\pm$0.48 91.67$\pm$0.36 82.30$\pm$0.53 87.90$\pm$0.87 89.58$\pm$0.26 81.48$\pm$2.41 86.83$\pm$0.87 45.83$\pm$0.85 42.36$\pm$1.77
 + TFB (Ours) 77.44$\pm$0.30 82.53$\pm$1.00 91.33$\pm$0.37 82.53$\pm$0.56 88.53$\pm$0.57 89.75$\pm$0.25 79.76$\pm$1.24 85.52$\pm$0.56 44.33$\pm$4.03 37.00$\pm$2.16
MAP - 76.90$\pm$0.97 81.08$\pm$2.48 91.61$\pm$0.44 82.59$\pm$0.28 85.73$\pm$0.19 90.09$\pm$0.28 79.98$\pm$0.87 86.58$\pm$0.79 43.40$\pm$4.98 38.54$\pm$3.40
 + TFB (Ours) 76.43$\pm$0.72 82.80$\pm$1.42 91.39$\pm$0.37 82.64$\pm$0.58 86.00$\pm$0.16 89.96$\pm$0.18 80.61$\pm$1.24 86.30$\pm$0.89 45.33$\pm$2.87 35.67$\pm$4.11
BLoB-Mean 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05 82.06$\pm$1.15 88.54$\pm$0.31 39.93$\pm$5.20 39.93$\pm$4.02
 + TFB (Ours) 77.81$\pm$0.36 83.33$\pm$0.19 91.76$\pm$0.48 83.81$\pm$0.39 87.80$\pm$0.16 90.11$\pm$0.28 82.93$\pm$1.54 87.64$\pm$0.51 39.67$\pm$7.32 37.33$\pm$6.65
ECE ($\downarrow$) MCD 16.13$\pm$0.54 13.69$\pm$1.11 6.73$\pm$0.71 13.05$\pm$0.99 9.76$\pm$0.71 7.95$\pm$0.17 13.63$\pm$1.18 9.27$\pm$0.60 30.91$\pm$3.57 33.08$\pm$1.40
ENS 14.72$\pm$0.17 13.45$\pm$1.19 6.59$\pm$0.45 11.17$\pm$0.92 8.17$\pm$0.86 7.35$\pm$0.55 11.37$\pm$1.82 7.21$\pm$1.13 18.92$\pm$6.03 26.80$\pm$3.23
LAP BP 4.18$\pm$0.11 9.26$\pm$3.08 5.27$\pm$0.51 3.50$\pm$0.78 8.93$\pm$0.34 1.93$\pm$0.22 7.83$\pm$1.49 7.80$\pm$1.99 14.49$\pm$0.57 13.17$\pm$2.14
BLoB 9.93$\pm$0.22 5.41$\pm$1.17 2.70$\pm$0.87 4.28$\pm$0.64 2.91$\pm$0.92 2.58$\pm$0.25 5.61$\pm$0.40 2.48$\pm$0.43 16.67$\pm$0.87 12.78$\pm$4.18
MLE - 17.02$\pm$0.46 16.35$\pm$0.68 7.00$\pm$0.53 13.83$\pm$0.65 9.77$\pm$0.81 8.69$\pm$0.21 14.45$\pm$2.19 10.78$\pm$0.50 32.46$\pm$2.60 38.41$\pm$4.44
 + TFB (Ours) 12.98$\pm$0.37 11.63$\pm$0.68 5.14$\pm$0.14 10.01$\pm$0.70 7.20$\pm$0.47 7.39$\pm$0.26 6.54$\pm$0.53 5.69$\pm$1.64 14.63$\pm$1.46 19.68$\pm$3.27
MAP - 18.71$\pm$0.74 15.77$\pm$1.60 6.62$\pm$0.64 14.26$\pm$0.92 12.19$\pm$0.55 8.40$\pm$0.25 16.46$\pm$0.44 11.36$\pm$0.58 34.79$\pm$3.76 38.50$\pm$2.18
 + TFB (Ours) 14.95$\pm$0.65 11.27$\pm$2.53 5.76$\pm$0.63 10.97$\pm$1.19 9.70$\pm$0.69 6.86$\pm$0.31 13.25$\pm$0.95 9.22$\pm$0.91 27.21$\pm$2.62 35.91$\pm$4.12
BLoB-Mean 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29 11.22$\pm$0.38 6.34$\pm$0.71 26.65$\pm$3.06 25.40$\pm$5.40
 + TFB (Ours) 8.16$\pm$0.48 6.48$\pm$0.36 2.44$\pm$0.50 3.83$\pm$0.43 2.67$\pm$0.18 3.10$\pm$0.59 6.69$\pm$1.63 3.61$\pm$0.87 18.45$\pm$6.75 20.53$\pm$6.27
NLL ($\downarrow$) MCD 0.83$\pm$0.01 0.99$\pm$0.10 0.45$\pm$0.06 0.64$\pm$0.03 0.62$\pm$0.08 0.49$\pm$0.01 1.03$\pm$0.02 0.61$\pm$0.03 1.91$\pm$0.18 2.02$\pm$0.15
ENS 0.75$\pm$0.02 0.80$\pm$0.11 0.38$\pm$0.03 0.55$\pm$0.02 0.45$\pm$0.05 0.42$\pm$0.05 0.72$\pm$0.07 0.44$\pm$0.03 1.40$\pm$0.18 1.50$\pm$0.13
LAP BP 0.56$\pm$0.00 1.18$\pm$0.02 1.04$\pm$0.01 0.51$\pm$0.00 0.94$\pm$0.00 0.43$\pm$0.00 1.17$\pm$0.01 1.11$\pm$0.00 1.27$\pm$0.01 1.28$\pm$0.00
BLoB 0.58$\pm$0.00 0.51$\pm$0.03 0.23$\pm$0.01 0.43$\pm$0.01 0.34$\pm$0.01 0.26$\pm$0.01 0.56$\pm$0.02 0.35$\pm$0.02 1.34$\pm$0.04 1.35$\pm$0.10
MLE - 0.88$\pm$0.04 1.20$\pm$0.11 0.46$\pm$0.04 0.68$\pm$0.01 0.61$\pm$0.06 0.52$\pm$0.01 1.07$\pm$0.06 0.72$\pm$0.06 1.91$\pm$0.16 2.25$\pm$0.21
 + TFB (Ours) 0.68$\pm$0.03 0.85$\pm$0.02 0.33$\pm$0.03 0.53$\pm$0.01 0.46$\pm$0.04 0.42$\pm$0.00 0.66$\pm$0.02 0.44$\pm$0.01 1.39$\pm$0.11 1.49$\pm$0.05
MAP - 0.99$\pm$0.07 1.12$\pm$0.23 0.46$\pm$0.03 0.74$\pm$0.07 0.79$\pm$0.02 0.52$\pm$0.01 1.19$\pm$0.04 0.83$\pm$0.06 1.97$\pm$0.13 2.32$\pm$0.10
 + TFB (Ours) 0.77$\pm$0.05 0.80$\pm$0.15 0.38$\pm$0.03 0.57$\pm$0.05 0.61$\pm$0.03 0.40$\pm$0.01 0.96$\pm$0.08 0.66$\pm$0.06 1.69$\pm$0.16 2.12$\pm$0.08
BLoB-Mean 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02 0.67$\pm$0.07 0.39$\pm$0.03 1.53$\pm$0.13 1.54$\pm$0.15
 + TFB (Ours) 0.55$\pm$0.01 0.53$\pm$0.04 0.23$\pm$0.02 0.40$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01 0.52$\pm$0.05 0.35$\pm$0.02 1.36$\pm$0.13 1.46$\pm$0.11
@@
For in-distribution experiments, we evaluate model performance on six commonsense reasoning tasks: Winogrande-Small (WG-S) and Winogrande-Medium (WG-M[50], ARC-Challenge (ARC-C) and ARC-Easy (ARC-E[11], Open Book Question Answering (OBQA[42], and BoolQ [10]. Furthermore, we use models fine-tuned on OBQA [42] to evaluate their generalization ability on out-of-distribution datasets: college-level chemistry (Chem) and physics (Phy) subsets of MMLU [20]. Label spaces and prompt templates are detailed in Appendix E.1.
@@
To assess uncertainty estimation, we measure Expected Calibration Error (ECE [45]) and Negative Log-Likelihood (NLL) on the test dataset. We also report Accuracy (ACC) to ensure models maintain strong performance. Additional evaluation details are provided in Appendix E.2.
Baselines. We compare TFB with state-of-the-art uncertainty estimation methods for LoRA-adapted LLMs, including ensemble-based method: Deep Ensemble (ENS[39, 4, 60], variational inference methods: Monte-Carlo Dropout (MCD[14], Bayesian LoRA by Backprop (BLoB[61], and post-training method: Laplace-LoRA (LAP[66]. For reference, we also include two standard PEFT baselines: Maximum Likelihood Estimation (MLE[23] and Maximum A Posteriori (MAP). All baselines are implemented following the protocols established in BLoB, detailed in Appendix E.4.
TFB Implementation.TFB can be directly applied to trained LoRA adapters without additional training. As indicated by the “TF?” column in Table 1, TFB is Training-Free and requires only LLM inference (), while the other methods need full retraining () or gradient estimation with Backpropagation (BP). We evaluate TFB on three off-the-shelf LoRA checkpoints: MLE, MAP, and the mean component of BLoB (obtained by discarding BLoB’s standard deviation matrix ${\bm{\Omega}}$). More details are included in Appendix E.4.
Table 1 shows results on comprehensive metrics for various methods applied to LoRA on Llama3.1-8B pre-trained weights. More empirical results on Llama2-7B can be found in Appendix F.10.
In-Distribution Results. The addition of TFB maintains competitive accuracy while substantially improving model calibration across in-distribution datasets. For ECE, TFB yields notable improvements when applied to different base methods: MLE+TFB reduces ECE to 5.14% on ARC-E (from 7.00%); similarly MAP+TFB and BLoB-Mean+TFB reduce ECE to 9.70% on OBQA (from 12.19%) and 3.83% on WG-M (from 9.37%), respectively. For NLL, TFB consistently produces better-calibrated predictions, with BLoB-Mean+TFB achieving strong performance across datasets: 0.23 on ARC-E (from 0.29), 0.33 on OBQA (from 0.37), and 0.27 on BoolQ (from 0.32). These improvements in both ECE and NLL demonstrate TFB’s effectiveness in enhancing model calibration while preserving accuracy on in-distribution tasks.
Out-of-Distribution Results. For out-of-distribution datasets, which represent a more challenging evaluation scenario, TFB continues to show benefits, though the performance gaps are generally smaller. In both Small Shift and Large Shift scenarios, TFB-enhanced methods maintain relatively strong performance, particularly in the Small Shift cases (ARC-C and ARC-E). However, there’s a noticeable performance drop in the Large Shift scenarios (Chem and Phy), which is expected given the significant domain difference. Even in these challenging cases, TFB-enhanced methods tend to maintain better calibration (lower ECE scores) compared to their base counterparts, suggesting improved reliability in out-of-distribution settings.
5.3 TFB Beyond the Low-Rank Isotropic Gaussians
@@
Table 2: Performance of TFB with different variational distribution families applied to BLoB-Mean on Llama3.1-8B pre-trained weights. FR: Full-rank isotropic Gaussian noises are applied to $\Delta{\bm{W}}$; C-STD: Standard deviation matrix ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$ is constant. The evaluation protocol strictly follows Table 1. “Rk.”: Average ranking of each method when compared to all other approaches on in-distribution datasets. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
ACC ($\uparrow$) BLoB-Mean 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05 2.50 82.06$\pm$1.15 88.54$\pm$0.31 39.93$\pm$5.20 39.93$\pm$4.02
 + TFB (FR) 75.57$\pm$0.25 83.20$\pm$0.65 91.58$\pm$0.67 82.19$\pm$1.09 88.73$\pm$0.41 89.46$\pm$0.17 2.83 81.33$\pm$0.82 88.06$\pm$0.75 42.00$\pm$2.16 41.33$\pm$5.44
 + TFB (C-STD) 76.35$\pm$0.08 83.20$\pm$0.33 91.33$\pm$0.70 81.79$\pm$0.51 88.20$\pm$0.57 89.65$\pm$0.08 3.00 81.73$\pm$0.68 88.18$\pm$0.65 43.00$\pm$1.41 39.33$\pm$3.86
 + TFB (Final) 77.81$\pm$0.36 83.33$\pm$0.19 91.76$\pm$0.48 83.81$\pm$0.39 87.80$\pm$0.16 90.11$\pm$0.28 1.67 82.93$\pm$1.54 87.64$\pm$0.51 39.67$\pm$7.32 37.33$\pm$6.65
ECE ($\downarrow$) BLoB-Mean 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29 4.00 11.22$\pm$0.38 6.34$\pm$0.71 26.65$\pm$3.06 25.40$\pm$5.40
 + TFB (FR) 10.42$\pm$0.29 7.45$\pm$0.88 2.01$\pm$1.03 4.36$\pm$0.68 3.70$\pm$1.04 3.62$\pm$0.10 2.67 7.19$\pm$1.40 3.29$\pm$1.03 17.78$\pm$1.01 19.14$\pm$4.01
 + TFB (C-STD) 9.23$\pm$0.20 5.98$\pm$0.32 2.94$\pm$0.67 3.86$\pm$0.45 3.17$\pm$0.21 2.82$\pm$0.62 1.83 6.89$\pm$0.89 2.76$\pm$0.88 18.27$\pm$2.52 19.45$\pm$3.46
 + TFB (Final) 8.16$\pm$0.48 6.48$\pm$0.36 2.44$\pm$0.50 3.83$\pm$0.43 2.67$\pm$0.18 3.10$\pm$0.59 1.50 6.69$\pm$1.63 3.61$\pm$0.87 18.45$\pm$6.75 20.53$\pm$6.27
NLL ($\downarrow$) BLoB-Mean 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02 3.67 0.67$\pm$0.07 0.39$\pm$0.03 1.53$\pm$0.13 1.54$\pm$0.15
 + TFB (FR) 0.60$\pm$0.01 0.53$\pm$0.03 0.23$\pm$0.02 0.43$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01 2.00 0.57$\pm$0.04 0.34$\pm$0.02 1.34$\pm$0.07 1.42$\pm$0.09
 + TFB (C-STD) 0.57$\pm$0.01 0.51$\pm$0.02 0.22$\pm$0.01 0.43$\pm$0.01 0.33$\pm$0.01 0.26$\pm$0.01 1.33 0.56$\pm$0.04 0.33$\pm$0.02 1.34$\pm$0.08 1.41$\pm$0.09
 + TFB (Final) 0.55$\pm$0.01 0.53$\pm$0.04 0.23$\pm$0.02 0.40$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01 1.50 0.52$\pm$0.05 0.35$\pm$0.02 1.36$\pm$0.13 1.46$\pm$0.11
@@
In this section, we consider two simple TFB variants with other families of Gaussians for modeling the variational distributions of ${\bm{W}}$: (i) Full-Rank Isotropic Gaussian (FR, ${\bm{\Sigma}}_{q}=\sigma_{q}^{2}{\bm{I}}$), and (ii) Constant Low-Rank Standard Deviation (C-STD, ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$). Similar to our final TFB, both distributions are controlled by a single $\sigma_{q}$ parameter and fit the maximal variance search in Eqn. 8. For fair comparison, we adopt the same optimal $\sigma_{q}^{*}$ search protocol as described in Sec. 5.1. Table 2 shows the performances of TFB and its variants applied to the mean of BLoB (more in Table 11 of Appendix F.6).
Table 3: Performance of different LLM backbones on the combined dataset.
@@
These results show that our final TFB outperforms both variants FR and C-STD across multiple metrics on in-distribution datasets, with notable improvements in calibration (ECE reduced by up to 15.77%) and accuracy (e.g., 77.81% on WG-S). While C-STD shows better NLL scores, the improvement comes at the cost of a significantly degraded overall performance—particularly in accuracy, where it performs the worst—making it impractical for real-world applications. Although our final TFB maintains strong performance on datasets with smaller distributional shifts, its advantages diminish on datasets with larger shifts in the domains of Physics and Chemistry.
Advantages of Final TFB’s Variational Low-Rank Isotropic Gaussians. Compared to TFB (FR) and TFB (C-STD), TFB (Final) offers additional advantages. It is computationally more efficient than FR with noise complexity of $O(rn)$ versus $O(mn)$. Furthermore, unlike C-STD whose variational distributions vary with different but equivalent LoRA matrix pairs (see Appendix F.6 for details), TFB (Final) produces consistent Bayesianization for all equivalent LoRAs satisfying ${\bm{B}}{\bm{A}}=\Delta{\bm{W}}$.
5.4 TFB Beyond the Llama3.1-8B Backbone
We conduct comprehensive experiments across multiple LLM backbones to validate our approach. Our experiments span several models from the Llama family [54, 12], including llama-2-7b-hf, Meta-Llama-3-8B, and Meta-Llama-3.1-8B. While we initially considered Llama-3.2-1B, we ultimately discarded its results due to poor adaptation performance with the smaller model architecture. We also extend our analysis to include Mistral-7B-v0.3 [26].
Following commonsense-170k [24, 59], we combine the 6 reasoning sub-tasks from the main experiments with one shared label space (Combined) and train the base LoRA adapters with MLE. Break-down statistics of each sub-dataset are available in Appendix F.7. Table 3 shows the results, demonstrating TFB’s effectiveness: it dramatically reduces ECE across all models (e.g., from 4.74% to 1.05% for Llama3.1-8B) while maintaining strong ACC and NLL scores.